# Coded Downlink Massive Random Access and a Finite de Finetti Theorem

### v1

@@ II-E4 Resource Allocation
Suppose that $\mathbf{X}=(X_{1},\cdots,X_{k})$ must satisfy $\sum_{i=1}^{k}X_{i}=r$, where $0\leq X_{i}\leq r$ takes on integer values. This can be thought of as a resource allocation problem where $r$ units of resources are allocated to $k$ randomly activated users (among a total of $n$ potential users). Consider the joint distribution induced by the process of allocating each unit of resource to the $k$ users uniformly at random one at a time. The resulting $(X_{1},\cdots,X_{k})$ is exchangeable, but $X_{i}$’s are not independent due to the constraint $\sum_{i=1}^{k}X_{i}=r$.
The source distribution in this setting is the multinomial distribution:
$\mathrm{Pr}(\mathbf{X}=\mathbf{x})=\begin{cases}\frac{r!}{x_{1}!\cdots x_{k}!}\left(\frac{1}{k}\right)^{r},&\text{if }\sum_{i=1}^{k}x_{i}=r\\ 0,&\text{otherwise}.\end{cases}$ (38)
The multinomial distribution is exchangeable. By Corollary 2, using the urn codebook achieves a rate of $H(\mathbf{X})$ bits with an overhead of $k\log(e)$ plus a logarithmic term.
@@
If we use the marginal distribution to construct the codebook, the codeword distribution would have been $q(\mathbf{x})=\prod_{i=1}^{k}p(x_{i})$, where $p(x_{i})$ is the marginal of the multinomial distribution. Then by Theorem 1, an achievable rate (up to an additional logarithmic term) is:
$\begin{split}R_{\mathrm{marginal}}&\approx H(\mathbf{X})+D(p\|q)\\ &=\sum_{x_{1},\cdots,x_{k}}p(\mathbf{x})\log\left(\frac{1}{\prod_{i=1}^{k}p(x_{i})}\right)\\ &=k\sum_{x_{1}}p(x_{1})\log\left(\frac{1}{p(x_{1})}\right)\\ &=kH(X_{1}).\end{split}$ (39)
The marginal distribution of the multinomial is a binomial distribution $p(x_{1})={{r}\choose{x_{1}}}\left(\frac{1}{k}\right)^{x_{1}}\left(1-\frac{1}{k}\right)^{r-x_{1}}$. The entropy of the marginal can be computed as
$H(X_{1})=\frac{1}{2}\log\left(2\pi e\left(\frac{r}{k}\right)\left(1-\frac{1}{k}\right)\ \right)+O\left(\frac{1}{r}\right).$ (40)
On the other hand, if we use an urn codebook, by Corollary 2 the following rate is achievable (to within an additional logarithmic term):
@@
$=$ $\frac{k-1}{2}\log(2\pi er)-\frac{k}{2}\log(k)+O\left(\frac{1}{r}\right)+k\log e.$
@@
Comparing $R_{\mathrm{marginal}}$ with $R_{\mathrm{urn}}$, we see that their difference is as follows:
$\begin{split}&R_{\mathrm{marginal}}-R_{\mathrm{urn}}\\ &\approx\frac{1}{2}\log(2\pi er)+\frac{k}{2}\log\left(1-\frac{1}{k}\right)-k\log e+O\left(\frac{1}{r}\right)\\ &\geq\frac{1}{2}\log(2\pi er)-\frac{\log(e)}{2}-k\log e+O\left(\frac{1}{r}\right)\\ &=\frac{1}{2}\log(2\pi r)-k\log e+O\left(\frac{1}{r}\right).\end{split}$ (42)
Therefore at fixed $k$, for large values of $r$, it is advantageous to use the urn codebook.
The value of $r$ for which the urn codebook outperforms a codebook constructed from the marginal distribution does not need to be as large as what (42) suggests. Consider the scenario where $r=k=2$. In this case, the entropy of the marginal can be computed directly as
$H(X_{1})=-\sum_{i=0}^{2}p_{i}\log p_{i}=1.5$ (43)
where $p_{i}={{2}\choose{i}}\left(\frac{1}{2}\right)^{2}$. Therefore $kH(X_{1})=3$.
On the other hand, the entropy of $H(X_{1},X_{2})=H(X_{1})=1.5$, since the value of $X_{2}$ is determined by the value of $X_{1}$. This gap is therefore $1.5$ bits. Now, the extra $k\log(e)$ cost comes from approximating the divergence between the multinomial distribution and the corresponding urn codebook. For finite $k$, the extra cost is bounded above by $\log\left(\frac{k^{k}}{k!}\right)=1<1.5$ for $k=2$. Therefore, the urn codebook is already better than the codebook constructed from the marginal for $r=k=2$.
The urn codebook in this case contains only two types of codewords: the $(1,\cdots,1)$ codeword and codewords with 00 and $2$ in 50% proportion each, whereas the codebook constructed from marginal distribution has codewords with mixed 0, 1, and 2 entries, which is not as effective.
@@ II-F Random Number of Active Users
So far in this paper, we have assumed that the number of active users $k$ is fixed and known in advance. This is not always a valid assumption, since user activities are typically random in a massive access scenario. In this section, we point out that if the number of active users is random, the previous results can be modified so they continue to hold in expectation.
Consider the scenario in which we wish to transmit i.i.d. sources to a set of randomly activated users. Let $K$ be a random variable that denotes the number of active users. The BS learns the activity pattern, thus the realization of $K$ in each transmission slot, then aim to transmit the sources $X_{1},\cdots,X_{K}$ to each of the $K$ active users, respectively. One way to do this is to first broadcast the realization of $K$ to all the active users using $H(K)$ bits, then use the results developed in earlier sections of this paper. However, we show here that this is not necessary, in case of i.i.d. sources. The following result is a counterpart of Corollary 1 for random number of active users, which shows that the minimum common message rate is essentially the expected entropy of the source.
@@
Fix the total number of users $n$. Let the number of randomly activated users $K\in[n]$ be a random variable. Let $a_{i}$ be the index of the $i$th active user, $i\in[K]$. Let $(X_{1},\cdots,X_{K})$ be i.i.d. $\sim p(x)$. The minimum common message rate $R^{*}$ for communicating each $X_{i}$ to its respective intended user $a_{i}$, $i=1,\cdots,K$, is bounded above as
$R^{*}\leq\mathbb{E}[K]H(p)+\log(\mathbb{E}[K]H(p)+1)+1.$ (44)
@@ III Finite de Finetti Theorem
In this section, we discuss the implications of Theorem 2 and Theorem 3 towards the study of the de Finetti theorem and finite exchangeable sequences. The original de Finetti theorem states that every infinitely extendable sequence of exchangeable random variables has a distribution that is equivalent to an i.i.d. mixture. This result is later generalized to finite exchangeable sequences in [5] where it is shown that if $(X_{1},\cdots,X_{k},\cdots,X_{d})$ is an exchangeable sequence of random variables, then the distribution of $(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ is close to being an i.i.d. mixture in the sense that
$\min_{q\in\mathcal{Q}}\mathrm{TV}(p,q)\leq\min\left\{\frac{|\mathcal{X}|k}{d},\frac{k(k-1)}{2d}\right\},$ (45)
where $\mathrm{TV}(p,q)=\sup_{\mathcal{X}^{k}}|p(\mathbf{x})-q(\mathbf{x})|$ and $\mathcal{Q}$ is the family of i.i.d. mixture distributions on $\mathcal{X}^{k}$.
The interesting part of this result is that the total variation distance can go to 00 even if $k,d\to\infty$. In the case where the alphabet size $|\mathcal{X}|$ is finite, the total variation distance goes to 00 as long as $\frac{k}{d}\to 0$. For general alphabets, $\frac{k^{2}}{d}\to 0$ is sufficient. Examples are presented in [5] to show that the scaling of (45) cannot be improved in general.
Although the total variation distance cannot be used to upper bound the relative entropy, they are related through Pinsker’s inequality [13] which states that
$\mathrm{TV}(p,q)\leq\sqrt{\frac{\ln(2)}{2}D(p\|q)}.$ (46)
Since the scaling in (45) is the best possible for general sequences, we have a natural converse for upper bounds on the relative entropy between i.i.d. mixtures and exchangeable and extendable sequences, i.e. the bounds cannot scale better than (45).
Recently, there has been an interest in establishing similar results as (45), but in terms of relative entropy. Given that $p(\mathbf{x})$ is the distribution of a $d$-extendable exchangeable sequence $(X_{1},\cdots,X_{k})$, we wish to upper bound
$\min_{q\in\mathcal{Q}}D(p\|q).$ (47)
It is shown in [11] through the use of information theoretic tools that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\frac{k(k-1)}{2(d-k-1)}\log|\mathcal{X}|,$ (48)
indicating that if the cardinality of the alphabet is finite, then $\frac{k^{2}}{d}\to 0$ is sufficient to ensure vanishing relative entropy between a $d$-extendable exchangeable distribution and the closest i.i.d. mixture distribution. As noted by the authors, this scaling is weaker than that of (45).
Given an exchangeable source $\mathbf{X}=(X_{1},\cdots,X_{k})$ with distribution $p(\mathbf{x})$, we show in Theorem 1 that the achievable rate is close to the entropy $H(\mathbf{X})$ plus a divergence term
@@
where $\mathcal{Q}$ is the family of i.i.d. mixture distributions on $\mathcal{X}^{k}$. Through the extended urn codebook construction, we show through Theorem 3 that for every exchangeable and $d$-extendable distribution $p(\mathbf{x})$, there exists an i.i.d. mixture distribution $q(\mathbf{x})$ such that
Thus, the extended urn codebook construction implies a relative entropy version of the de Finetti theorem.
@@
Let $(X_{1},\cdots,X_{k},\cdots,X_{d})$ be a sequence of exchangeable random variables taking values in $\mathcal{X}^{d}$. Let $p(x_{1},\cdots,x_{k})$ denote the distribution of $(X_{1},\cdots,X_{k})$. Then,
$\begin{split}\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\bigg{\{}(|\mathcal{X}|-&1)\log\left(\frac{d-1}{d-k}\right),\\ &-\log\left(1-\frac{k(k-1)}{2d}\right)\bigg{\}},\end{split}$ (51)
@@
The proof follows directly from Theorem 3 and the following useful inequality from [14]:
$\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)\leq-\log\left(1-\frac{k(k-1)}{2d}\right).$ (52)
Clearly, the scaling of $d$ and $k$ required to drive $D(p\|q)$ to zero in Theorem 5 matches that of (45), and is therefore the best possible for general exchangeable sequences due to (46) and the aforementioned reasoning.
In fact, not only the scaling cannot be improved, the first term in the bound (50) itself, i.e.,
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)$ (53)
is also the best possible upper bound for general exchangeable and $d$-extendable sequences. We establish this fact by showing that for a particular exchangeable and $d$-extendable distribution $h$, the above upper bound is achieved.
@@
Let $d>k$ and $h(x_{1},\cdots,x_{k})$ be the distribution of $k$ draws without replacement from an urn containing elements $\{1,2,\cdots d\}$, i.e., $|\mathcal{X}|=d$. Then, for every i.i.d. mixture distribution $q(x_{1},\cdots,x_{k})$, we have that
$D(h\|q)\geq\log\left(\frac{d^{k}}{d^{\underline{k}}}\right).$ (54)
@@
The proof takes the idea from [5] and is deferred to Appendix G. ∎
Thus, for this particular $h(x_{1},\cdots,x_{k})$, the minimum $D(h\|q)$ over $q\in\mathcal{Q}$ is precisely the right-hand side of (53) and (54). For $d=k$, this lower bound is $\log\left(\frac{k^{k}}{k!}\right)\approx k\log e$.
Note that in this case, we have $|\mathcal{X}|=d$ for which the second term in the minimum in (50) is always larger than the first term. For general exchangeable and $d$-extendable $p(x_{1},\cdots,x_{k})$, e.g., with smaller $|\mathcal{X}|$, the minimum $D(p\|q)$ may be smaller than the right-hand side of (53).

### v4

@@ II-E4 Resource Allocation
Suppose that $\mathbf{X}=(X_{1},\cdots,X_{k})$ must satisfy $\sum_{i=1}^{k}X_{i}=\alpha$, where $\alpha\in\mathbb{N}$ and $0\leq X_{i}\leq\alpha$ takes on integer values. This can be thought of as a resource allocation problem where $\alpha$ units of resources are allocated to $k$ randomly activated users (among a total of $n$ potential users). Consider the joint distribution induced by the process of allocating each unit of resource to the $k$ users uniformly at random one at a time. The resulting $(X_{1},\cdots,X_{k})$ is exchangeable, but $X_{i}$’s are not independent due to the constraint $\sum_{i=1}^{k}X_{i}=\alpha$.
The distribution of the sources in this setting is the multinomial distribution:
$\mathrm{Pr}(\mathbf{X}=\mathbf{x})=\begin{cases}\frac{\alpha!}{x_{1}!\cdots x_{k}!}\left(\frac{1}{k}\right)^{\alpha},&\text{if }\sum_{i=1}^{k}x_{i}=\alpha\\ 0,&\text{otherwise}.\end{cases}$ (57)
The multinomial distribution is exchangeable. By Corollary 2, using the urn codebook achieves a common message length of $H(\mathbf{X})$ bits with an overhead of $k\log(e)$ plus a logarithmic term.
@@
If we use the marginal distribution to construct the codebook, the codeword distribution would have been $q(\mathbf{x})=\prod_{i=1}^{k}p(x_{i})$, where $p(x_{i})$ is the marginal of the multinomial distribution. Then by Theorem 1, an achievable common message length (up to an additional logarithmic term) is
$R_{\mathrm{marginal}}$$\approx$ $H(\mathbf{X})+D(p\|q)$ (58)
$=$ $\sum_{x_{1},\cdots,x_{k}}p(\mathbf{x})\log\left(\frac{1}{\prod_{i=1}^{k}p(x_{i})}\right)$ (59)
$=$ $k\sum_{x_{1}}p(x_{1})\log\left(\frac{1}{p(x_{1})}\right)$ (60)
$=$ $kH(X_{1}).$ (61)
The marginal distribution of the multinomial is a binomial distribution $p(x_{1})={{\alpha}\choose{x_{1}}}\left(\frac{1}{k}\right)^{x_{1}}\left(1-\frac{1}{k}\right)^{\alpha-x_{1}}$. The entropy of the marginal can be computed as
$H(X_{1})=\frac{1}{2}\log\left(2\pi e\left(\frac{\alpha}{k}\right)\left(1-\frac{1}{k}\right)\ \right)+O\left(\frac{1}{\alpha}\right).$ (62)
On the other hand, if we use an urn codebook, by Corollary 2 the following common message length is achievable (to within an additional logarithmic term):
@@
$=\frac{k-1}{2}\log(2\pi e\alpha)-\frac{k}{2}\log(k)+O\left(\frac{1}{\alpha}\right)+k\log e,$ (63)
@@
Comparing $R_{\mathrm{marginal}}$ with $R_{\mathrm{urn}}$, we see that their difference can be lower bounded as follows:
$R_{\mathrm{marginal}}-R_{\mathrm{urn}}$
$\approx\frac{1}{2}\log(2\pi e\alpha)+\frac{k}{2}\log\left(1-\frac{1}{k}\right)-k\log e+O\left(\frac{1}{\alpha}\right)$ (64)
$\geq\frac{1}{2}\log(2\pi e\alpha)-\frac{\log(e)}{2}-k\log e+O\left(\frac{1}{\alpha}\right)$ (65)
$=\frac{1}{2}\log(2\pi\alpha)-k\log e+O\left(\frac{1}{\alpha}\right).$ (66)
Therefore at fixed $k$, for large values of $\alpha$, it is advantageous to use the urn codebook.
The value of $\alpha$ for which the urn codebook outperforms a codebook constructed from the marginal distribution does not need to be as large as what (66) suggests. Consider the scenario where $\alpha=k=2$. In this case, the entropy of the marginal can be computed directly as
$H(X_{1})=-\sum_{i=0}^{2}p_{i}\log p_{i}=1.5,$ (67)
where $p_{i}={{2}\choose{i}}\left(\frac{1}{2}\right)^{2}$. Therefore, $kH(X_{1})=3$.
On the other hand, the entropy of $H(X_{1},X_{2})=H(X_{1})=1.5$, since the value of $X_{2}$ is determined by the value of $X_{1}$. This gap is therefore $1.5$ bits. Now, the extra $k\log(e)$ cost comes from approximating the divergence between the multinomial distribution and the corresponding urn codebook. For finite $k$, the extra cost is bounded from above by $\log\left(\frac{k^{k}}{k!}\right)=1<1.5$ for $k=2$. Therefore, the urn codebook is already better than the codebook constructed from the marginal for $\alpha=k=2$.
The urn codebook in this case contains only two types of codewords: $(1,\cdots,1)$ codeword and codewords with 00 and $2$ in 50% proportion each, whereas the codebook constructed from marginal distribution has codewords with mixed 0, 1, and 2 entries, which is not as effective.
@@ II-F Random Number of Active Users
So far in this paper, we have assumed that the number of active users $k$ is fixed and known in advance. This is not always realistic, because user activities are typically random in a massive access scenario. In this section, we consider a scenario in which we wish to transmit sources $(X_{1},\cdots,X_{K})$ to a set of randomly activated users. Let $K$ be a random variable which denotes the number of active users. The BS learns the activity pattern, thus the realization of $K$ in each transmission slot, then aims to transmit the sources $(X_{1},\cdots,X_{K})$ to each of the $K$ active users, respectively.
For the problem of communicating general exchangeable sources of variable size $K$, because most random vectors do not have well defined distributions when the size of the vector can vary, devising a general coding strategy for such sources is challenging. However, what one can do is to construct a library of codebooks, one for each possible value of $K$. After the set of active users are detected, the BS first broadcasts the actual number of active users using $H(K)$ bits, then uses the codebook designed for the sources $(X_{1},\cdots,X_{K})$ to communicate to the $K$ users. In this way, the total expected common message length is simply the average common message length averaged over $K$ plus $H(K)$ bits.
However, in the special case that the sources are i.i.d., it turns out that it is not necessary to communicate the number of active users, thus avoiding the $H(K)$ bits overhead. The following result is a counterpart of Corollary 1 for random number of active users, which shows that the minimum common message length is essentially the expected entropy of the sources.
@@
Fix the total number of users $n$. Let the number of active users $K\in[n]$ be a random variable. Let $a_{i}$ be the index of the $i$-th active user, $i\in[K]$. Let $(X_{1},\cdots,X_{K})$ be i.i.d. $\sim p(x)$. The minimum common message length $R^{*}$ for communicating each $X_{i}$ to its respective intended user $a_{i}$, $i=1,\cdots,K$, is bounded from above as
$R^{*}<\mathbb{E}[K]H(p)+\log(\mathbb{E}[K]H(p)+1)+2.$ (68)
@@ III Finite de Finetti Theorem
In this section, we discuss the connection between the codebook construction for downlink massive random access and the study of finite exchangeable sequences and the finite de Finetti theorems.
III-A Existing Finite de Finetti Theorems
The original de Finetti theorem states that every infinitely extendable sequence of exchangeable random variables has a distribution that is equivalent to an i.i.d. mixture. This result can be generalized to finite exchangeable sequences in what is known as the finite de Finetti theorem. Specifically, it is shown in [5] that if $(X_{1},\cdots,X_{k},\cdots,X_{d})$ is an exchangeable sequence of random variables, then the distribution of $(X_{1},\cdots,X_{k})\sim p(\mathbf{x})$ is close to being an i.i.d. mixture in the sense that
$\min_{q\in\mathcal{Q}}\mathrm{TV}(p,q)\leq\min\left\{\frac{k(k-1)}{2d},\frac{|\mathcal{X}|k}{d}\right\},$ (69)
where $\mathrm{TV}(p,q)=\frac{1}{2}\sum_{\mathbf{x}=\mathcal{X}^{k}}|p(\mathbf{x})-q(\mathbf{x})|$ and $\mathcal{Q}$ is the family of i.i.d. mixture distributions on $\mathcal{X}^{k}$. This is an upper bound on how close finite exchangeable sequences are to being i.i.d. mixture in terms of TV distance.
An interesting implication of (69) is that the TV distance between the distribution of a $d$-extendable exchangeable sequence $(X_{1},\cdots,X_{k})$ and the closest i.i.d. mixture distribution can go to 00 even when $k\to\infty$, if $d$ also goes to infinity in certain ways. In the case of finite alphabet size, i.e., $|\mathcal{X}|<\infty$, the TV distance goes to 00 as long as $\frac{k}{d}\to 0$, and for general alphabets as long as $\frac{k^{2}}{d}\to 0$. Examples are presented in [5] that show these scalings cannot be improved upon in general.
There has been a recent interest in establishing similar results as (69), but in terms of KL divergence. Given that $p(\mathbf{x})$ is the distribution of a $d$-extendable exchangeable sequence $\mathbf{X}=(X_{1},\cdots,X_{k})$, the aim is to upper bound
@@ III-A Existing Finite de Finetti Theorems
Two upper bounds on (70) have recently been developed in [13] and [14]. The first bound is independent of the alphabet size, but depends on the source distribution. It states that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\frac{k(k-1)}{2(d-k+1)}H(X_{1}).$ (71)
The second bound, which depends on the alphabet size, states that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq(|\mathcal{X}|-1)\frac{k(k-1)}{2(d-1)(d-k+1)}.$ (72)
For the important special case where the exchangeable sequence is not extendable, i.e., $d=k$, the two upper bounds taken together imply that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\left\{\frac{k(k-1)H(X_{1})}{2},\frac{(|\mathcal{X}|-1)k}{2}\right\}.$ (73)
We see that the first argument of the min operation depends on $k$ as $O(k^{2})$, while the second argument depends on $|\mathcal{X}|$ and $k$ as $O(|\mathcal{X}|k)$. In this paper, we show that both of these scalings can be improved.
For the case of $d>k$, (71) and (72) imply that the KL divergence between an exchangeable sequence and its best i.i.d. mixture approximation goes to 00 as long as $\frac{k}{d}\to 0$ for the finite alphabet size case, and for general alphabets as long as $\frac{k^{2}}{d}\to 0$. These are the same scalings as implied by (69) for TV distance.
These scalings are optimal not only for TV distance but also for KL divergence. This is because although the TV distance cannot be used to upper bound KL divergence, the reverse is possible through Pinkser’s inequality [20]. Pinsker’s inequality states that
$\mathrm{TV}(p,q)\leq\sqrt{\frac{1}{2\log(e)}D(p\|q)}.$ (74)
Since the scalings in (69) are the best possible for TV distance for general exchangeable sequences, we have a natural converse for upper bounding the KL divergence between the i.i.d. mixtures and the exchangeable and extendable sequences, i.e. the bounds cannot scale better than (69).
Although the scalings of upper bounds (71) and (72) match the optimal scalings of [5], the bound (71) is distribution dependent as it contains the entropy term $H(X_{1})$. Using the approach developed in this paper, we show that it is possible to remove this distribution dependency.
III-B An Improved Finite de Finetti Theorem
In the context of downlink massive random access, the minimum divergence (70) is essentially the overhead for the minimum achievable common message length beyond the joint entropy $H(\mathbf{X})$, according to Theorem 1. For this reason, the characterizations of the common message length for the urn codebook construction in Theorem 2 and for the extended urn codebook construction in Theorem 3 immediately give us the following finite de Finetti theorem based on KL divergence.
@@ III-B An Improved Finite de Finetti Theorem
Let $(X_{1},\cdots,X_{k},\cdots,X_{d})$ be a sequence of exchangeable random variables taking values in $\mathcal{X}^{d}$, where $\mathcal{X}$ is a finite or countable alphabet. Let $p(x_{1},\cdots,x_{k})$ denote the distribution of $(X_{1},\cdots,X_{k})$. The minimum divergence is bounded as
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\log\left(\frac{d^{k}}{d^{\underline{k}}}\right).$ (75)
For the case of $d=k$, the right-hand side of the above is further upper bounded as $\log\left(\frac{k^{k}}{k!}\right)<k\log(e)$.
When the alphabet size is finite, we also have that
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\begin{cases}|\mathcal{X}|\log(k+1)&\mathrm{if}\ d=k\\ \displaystyle(|\mathcal{X}|-1)\log\left(\frac{d-1}{d-k}\right)&\mathrm{if}\ d>k.\end{cases}$ (76)
@@
The proof follows directly from Theorem 2 and Theorem 3. ∎
For the special case where the exchangeable sequence is not extendable, i.e., $d=k$, this finite de Finetti bounds is a consequence of Theorem 2, i.e.,
$\min_{q\in\mathcal{Q}}D(p\|q)\leq\min\left\{k\log(e),|\mathcal{X}|\log(k+1)\right\}.$ (77)
Comparing to (73), we see that the new bound improves the scaling for the finite alphabet size case from $O(|\mathcal{X}|k)$ to $O(|\mathcal{X}|\log(k))$, and for the general alphabet case from $O(k^{2})$ to $O(k)$. Furthermore, the first term of (77) is distribution independent, in contrast to the first term in (73) which depends on the entropy $H(X_{1})$.
For the case of $d>k$, we first analyze the scaling of (75). By using the following useful inequality from [21], which states that for $d>\frac{1}{2}k(k-1)$,
$\log\left(\frac{d^{k}}{d^{\underline{k}}}\right)\leq-\log\left(1-\frac{k(k-1)}{2d}\right),$ (78)
we see that the scaling of $d$ and $k$ required to drive the minimum $D(p\|q)$ to zero in the bound (75) is that $\frac{k^{2}}{d}\to 0$. For the finite alphabet size case, the scaling required in the bound (76) is that $\frac{k}{d}\to 0$. These scaling results match that of (71) and (72), and also that of (69). These are therefore the best possible scaling for reasons due to Pinsker’s inequality, as mentioned before.
A key feature of the bound (75) as compared to (71) is that it is distribution independent. On the other hand, the second case (for $d>k$) in (76) can be larger or smaller than (72), as analyzed in [10]. The second case in (76) is a result of [10, Theorem 4.5], which bounds the divergence between sampling with and without replacement from an urn containing elements from set $\mathcal{X}$. The bound (72) is derived using a similar argument, but uses an alternative bound from [9] instead of [10, Theorem 4.5]. A detailed comparison of these two bounds can be found in [10].
We conclude this section by pointing out that the bound (75) is not only scaling optimal, but also the best possible upper bound for general exchangeable and $d$-extendable sequences. We establish this fact by showing that for a particular exchangeable and $d$-extendable distribution $r$, the above upper bound is achieved.
@@
Fix $d\geq k$. Let $r(x_{1},\cdots,x_{k})$ be the distribution of $k$ draws without replacement from an urn containing elements $\{1,2,\cdots d\}$, i.e., $|\mathcal{X}|=d$. Then, for every i.i.d. mixture distribution $q(x_{1},\cdots,x_{k})$, we have that
$D(r\|q)\geq\log\left(\frac{d^{k}}{d^{\underline{k}}}\right).$ (79)
@@
The proof takes inspiration from [5] and is deferred to Appendix G. ∎
Thus, for this particular $r(x_{1},\cdots,x_{k})$, defined when $d\geq k$, the minimum $D(r\|q)$ over $q\in\mathcal{Q}$ is precisely the right-hand side of (75) and (79). In the case of $d=k$, this lower bound is $\log\left(\frac{k^{k}}{k!}\right)\approx k\log e$.
Note that in this example, we have $|\mathcal{X}|=d$, for which the right-hand side of (75) is always larger than the right-hand side of (76). For general exchangeable and $d$-extendable $p(x_{1},\cdots,x_{k})$, e.g., with smaller $|\mathcal{X}|$, the minimum $D(p\|q)$ may be smaller than the right-hand side of (75).